# Gaussian Splatting with NeRF-based Color and Opacity

### v2

@@ 3 Viewing Direction Gaussian Splatting
Chair Drums Lego Mic Materials Ship Hotdog Ficus Avg.
@@
Gaussian Splatting
GS algorithm creates the radiance field representation by a sequence of optimization steps of 3D Gaussian parameters (i.e., position, covariance, opacity, and SH colors). The key to the efficiency of GS is the rendering process, which uses projections of Gaussian components.
In GS, we use a dense set of 3D Gaussians
where $\mathrm{m}_{i}$ is position, $\Sigma_{i}$ is covariance, $\sigma_{i}$ is opacity, and $c$ is SH colors of $i$-th component.
GS starts with a limited number of points and then employs a strategy that involves creating new components and eliminating unnecessary ones. Every hundred iterations, GS eliminates any Gaussians with an opacity lower than a specific limit. At the same time, new Gaussians are generated in unoccupied areas of 3D space to fill vacant areas. The optimization process in GS is complicated, but it can be trained very effectively thanks to robust implementation and the usage of CUDA kernels.
@@
NVDiffRec (Munkberg et al., 2022) 27.98 21.77 26.97 40.44 24.31 30.74 28.70
NVDiffMC (Hasselgren et al., 2022) 25.93 30.85 26.27 38.44 22.18 29.60 28.88
Ref-NeRF (Verbin et al., 2022) 30.41 29.14 29.92 45.19 25.29 33.99 32.32
NeRO (Liu et al., 2023) 25.53 30.26 29.20 38.70 26.46 28.89 29.84
ENVIDR (Liang et al., 2023) 28.46 38.89 32.73 41.59 26.11 29.48 32.88
Guassian Splatting (Kerbl et al., 2023) 27.24 27.69 28.32 45.68 20.99 32.32 30.37
GaussianShader (Jiang et al., 2023) 27.90 30.98 28.32 45.86 26.21 32.39 31.94
VDGS (Our) 26.99 27.87 28.18 45.39 21.53 32.16 30.35
@@
NVDiffRec (Munkberg et al., 2022) 0.963 0.858 0.951 0.996 0.928 0.973 0.945
NVDiffMC (Hasselgren et al., 2022) 0.940 0.940 0.940 0.995 0.886 0.965 0.944
Ref-NeRF (Verbin et al., 2022) 0.949 0.956 0.955 0.995 0.910 0.972 0.956
NeRO (Liu et al., 2023) 0.949 0.974 0.971 0.995 0.929 0.956 0.962
ENVIDR (Liang et al., 2023) 0.961 0.991 0.980 0.996 0.939 0.949 0.969
Guassian Splatting (Kerbl et al., 2023) 0.930 0.937 0.951 0.996 0.895 0.971 0.947
GaussianShader (Jiang et al., 2023) 0.931 0.965 0.950 0.996 0.929 0.971 0.957
VDGS (Our) 0.938 0.944 0.949 0.996 0.899 0.971 0.949
@@
NVDiffRec (Munkberg et al., 2022) 0.045 0.297 0.118 0.011 0.169 0.076 0.119
NVDiffMC (Hasselgren et al., 2022) 0.077 0.312 0.157 0.014 0.225 0.097 0.147
Ref-NeRF (Verbin et al., 2022) 0.051 0.307 0.087 0.013 0.118 0.082 0.109
NeRO (Liu et al., 2023) 0.074 0.094 0.050 0.012 0.089 0.110 0.072
ENVIDR (Liang et al., 2023) 0.049 0.067 0.051 0.011 0.116 0.139 0.072
Guassian Splatting (Kerbl et al., 2023) 0.047 0.161 0.079 0.007 0.126 0.078 0.083
GaussianShader (Jiang et al., 2023) 0.045 0.121 0.076 0.007 0.079 0.078 0.068
VDGS (Our) 0.049 0.143 0.073 0.007 0.101 0.079 0.075
@@
Table 3: The quantitative comparisons (PSNR / SSIM / LPIPS) on Shiny Blender dataset (Verbin et al., 2022).
Viewing Direction Gaussian Splatting
Viewing Direction Gaussian Splatting
In our approach, we use GS to model 3D shapes, while a NeRF-based neural network produces colors and opacity. Our model consists of the Gaussian Splatting representation:
and the MLP network:
$\mathcal{F}_{VDGS{}}(\mathrm{m},\Sigma,c,{\bf d};\Theta)=\Delta\sigma({\bf d}),$
which takes parameters of Gaussian distribution and viewing direction ${\bf d}$ to returns emitted color update ${\bf\Delta c}({\bf d})$ and volume density update $\Delta\sigma({\bf d})$. The model is parameterized by $\bf\Theta$ and trained to change color and opacity depending on viewing direction.
The final Viewing Direction Gaussian Splatting model is given by:
where
$\Delta\sigma_{i}({\bf d})=\mathcal{F}_{VDGS{}}(\mathrm{m}_{i},\Sigma_{i},c_{i},{\bf d}).$
In practice, we have two components: GS and NeRF. GS is used to produce the shape and color of the objects. All Gaussian components have color and transparent, which do not depend on viewing direction. Thanks to GS, we model the universal colors of 3D objects. Depending on the camera position, the NeRF component can add minor color changes and transparency to the objects. Our model can add viewing direction-based changes. Therefore, it allows the modeling of light reflection and transparency of the objects.
4 Experiments
Datasets
To comprehensively validate the effectiveness of VDGS, we conducted an evaluation on various datasets: a) widely-used NVS dataset, i.e., the NeRF Synthetic (Mildenhall et al., 2020), b) real-world, large-scale scenes dataset, i.e., the Tanks and Temples (Knapitsch et al., 2017), and c) reflective objects datasets, i.e., the Shiny Blender (Verbin et al., 2022).
Dataset: Tanks and Temples
We also evaluated our model on larger-scale environments by comparing our solution on the Tanks and Temples dataset (Knapitsch et al., 2017). Tab. 1 and visual comparisons in Fig. 3 shows that we are able to better model light reflexes and shadows than either GS and NeRF.
Dataset: NeRF Synthetic
We first evaluate our model on the NeRF Synthetic dataset (Mildenhall et al., 2020), which contains objects with complex geometry and realistic non-Lambertian materials. We show the quantitative results in Tab. 2 and visual comparisons in Fig. 2. The results show that our approach achieves better results than both Gaussian Splatting and neural rendering methods.
Dataset: Shiny Blender
Viewing Direction Gaussian Splatting can plausible model shadows, light reflection, and transparency in 3D objects. Therefore, we decided to evaluate our hybrid approach on the Shiny Blender dataset (Verbin et al., 2022). Once again, visual comparisons in Fig. 4 and quantitative results in Tab. 3 show that we obtain comparable results to NeRF and GS approaches but slightly worse than dedicated method GaussianShader (Jiang et al., 2023).
5 Conclusions
@@
This paper presents a new neural rendering strategy that leverages two main concepts, NeRF and GS. We represent a 3D scene with a set of Gaussian components and a neural network that can change the color and opacity of Gaussians concerning viewing direction. This approach inherited the best elements from the NeRF and GS methods. We have quick training and inference time and we are able to model shadows, light reflections, and transparency of 3D objects.
The experimental section shows that Viewing Direction Gaussian Splatting gives better results than NeRF as well as the Gaussian Splatting model.

### v4

@@ 4 Experiments
Chair Drums Lego Mic Mat. Ship Hot. Ficus Avg.
@@
NVDiffRec 27.98 21.77 26.97 40.44 24.31 30.74 28.70
NVDiffMC 25.93 30.85 26.27 38.44 22.18 29.60 28.88
Ref-NeRF 30.41 29.14 29.92 45.19 25.29 33.99 32.32
NeRO 25.53 30.26 29.20 38.70 26.46 28.89 29.84
ENVIDR 28.46 38.89 32.73 41.59 26.11 29.48 32.88
GS 27.24 27.67 28.28 45.70 21.01 32.30 30.37
GaussianShader 27.90 30.98 28.32 45.86 26.21 32.39 31.94
VDGS (Our) 26.86 28.54 28.56 45.84 21.71 32.19 30.62
@@
NVDiffRec 0.963 0.858 0.951 0.996 0.928 0.973 0.945
NVDiffMC 0.940 0.940 0.940 0.995 0.886 0.965 0.944
Ref-NeRF 0.949 0.956 0.955 0.995 0.910 0.972 0.956
NeRO 0.949 0.974 0.971 0.995 0.929 0.956 0.962
ENVIDR 0.961 0.991 0.980 0.996 0.939 0.949 0.969
GS 0.931 0.937 0.951 0.997 0.895 0.972 0.947
GaussianShader 0.931 0.965 0.950 0.996 0.929 0.971 0.957
VDGS (Our) 0.925 0.944 0.950 0.997 0.896 0.971 0.947
@@
NVDiffRec 0.045 0.297 0.118 0.011 0.169 0.076 0.119
NVDiffMC 0.077 0.312 0.157 0.014 0.225 0.097 0.147
Ref-NeRF 0.051 0.307 0.087 0.013 0.118 0.082 0.109
NeRO 0.074 0.094 0.050 0.012 0.089 0.110 0.072
ENVIDR 0.049 0.067 0.051 0.011 0.116 0.139 0.072
GS 0.048 0.162 0.080 0.007 0.126 0.078 0.083
GaussianShader 0.045 0.121 0.076 0.007 0.079 0.078 0.068
VDGS (Our) 0.050 0.142 0.072 0.007 0.107 0.078 0.076
@@
Table 3: Quantitative comparisons (PSNR / SSIM / LPIPS) in the Shiny Blender dataset (Verbin et al., 2022) with NVDiffRec (Munkberg et al., 2022) NVDiffMC (Hasselgren et al., 2022) Ref-NeRF (Verbin et al., 2022) NeRO (Liu et al., 2023) ENVIDR (Liang et al., 2023) Guassian Splatting (Kerbl et al., 2023) GaussianShader (Jiang et al., 2023).
Car Ball Helmet Teapot Toaster Coffee Avg.
GS 27.24 27.67 28.28 45.70 21.01 32.30 30.37
VDGS (Our) 26.86 28.54 28.56 45.84 21.71 32.19 30.62
Pre-VDGS (Our) 27.29 27.81 28.37 45.71 21.17 32.33 30.45
Table 4: Comparison between original GS model 30k epochs, VDGS 30k epochs and pre-trained GS standard model with additional 20k epochs with VDGS on Shiny Blender (Verbin et al., 2022) data. We called the pre-trained model as Pre-VDGS.
NeRF Synthetic PSNR $\uparrow$
VDGS Chair Drums Lego Mic Materials Ship Hotdog Ficus Avg.
C$+$ 35.87 26.31 35.71 35.27 30.02 27.66 37.77 35.08 32.96
C* 35.91 26.20 35.73 35.44 29.96 29.63 37.76 35.03 33.21
O$+$ 35.60 26.14 34.41 35.33 29.97 28.70 37.09 34.93 32.77
O* 35.97 26.17 35.40 34.76 30.67 30.94 38.04 34.98 33.37
OC$+$ 35.77 26.28 31.17 33.70 29.49 17.91 36.68 35.24 30.78
OC* 36.21 26.03 35.44 34.59 30.66 30.88 38.03 34.29 33.27
Table 5: Ablation study of different versions of VDGS. We compare different models: Color Add (C$+$), Color Multiply C*, Opacity Add (O$+$), Opacity Multiply (O*) Opacity and Color Add (OC+), Opacity and Color Multiply (OC*). As we can see, different versions considered have good scores in various data sets and measures. In our paper, we chose VDGS Opacity Multiply (Our) as a final model since we obtained the largest number of best scores on the PSNR measure.
Pre-trained GS
Normally, our model simultaneously trains GS and NeRF components. In Table 4, we present results on the pre-trained classical GS model. We use a fully trained GS model, and then we additionally train our neural network. As can be seen, such an approach gives better results than the original GS. Furthermore, we obtain the best scores when we train both components simultaneously.
Implementation Details
We implement VDGS based on the original GS repository using the PyTorch and TinyCUDA frameworks. We trained our model on a single NVIDIA RTX 3090 GPU. We also utilized a 3D structured hash grid with 12 levels of 3D embedding ranging from 16 to 512 resolutions. The maximum hash table size was $2^{14}$ with a feature dimension of 2. For our NeRF-based network, we used a 2-layer MLP with LeakyReLU activation. We made sure to replicate the same configuration for all scenes.
Training and inference time
In contrast to GS in VDGS, we use a viewing direction neural network. Therefore, we have slightly longer training and inference time, as shown in Table 6. However, in practice, our model can work in real-time during inference.
Ablation Study
Our VDGS updates only opacity. It is a natural question of how our model works when also color changes. In the ablation study, we verified different versions of Viewing Direction Gaussian Splatting. In Table 5, we list the PSNR values for a different version of our model. At the same time, in the appendix, we present the full results of PSNR/SSIM/LPIPS measures.
As we can see, different versions of Viewing Direction Gaussian Splatting result in high scores for various data sets. In our paper, we chose VDGS Opacity Multiply as a final model since we obtained the highest number of best scores on the PSNR measure on the data containing real scenes. It should also be highlighted that the model VDGS Color Multiply gets the best scores on the NeRF Synthetic dataset. Models that use both color and opacity also obtain good results but are slightly worse than models that use only one of such elements. We believe this effect may be attributed to the dependency of opacity on color. Therefore, joint training is not effective.
Mip-NeRF360 Tanks&Temples Deep Blending
Train FPS Train FPS Train FPS
Plenoxels 25m49s 6.79 25m5s 13.0 27m49s 11.2
INGP-Base 5m37s 11.7 5m26s 17.1 6m31s 3.26
INGP-Big 7m30s 9.43 6m59s 14.4 8m 2.79
M-NeRF360 48h 0.06 48h 0.14 48h 0.09
GS-30K 41m33s 134 26m54s 154 36m2s 137
GS-30K* 26m21s 89.28 14m50s 94.07 24m04s 87.95
VDGS-30K* (Our) 46m11s 41.35 29m45s 38.53 41m29s 44.72
Table 6: Comparison of training time and frame rate. We train VDGS on the RTX 3090 GPU. We reproduce the result of GS on the RTX 3090 GPU. VDGS has slightly slower training and inference time compared to GS. However, it still yields real-time FPS superior to previous NeRF-based methods.
5 Limitations
Our model can outperform the original GS at multiple tasks, such as rendering glass surfaces and shadows or eliminating artifacts. However, due to its generality, it cannot reproduce the results of other solutions specifically designed to tackle selected niche problems, like accurate light reflections, by better simulating physical attributes.
6 Discussion and Conclusions
@@
This paper presents a new neural rendering strategy that leverages two main concepts, NeRF and GS. We represent a 3D scene with a set of Gaussian components and a neural network that can change the color and opacity of Gaussians concerning viewing direction. This approach inherited the best elements from the NeRF and GS methods. We observed rapid training and inference, and we can model shadows, light reflections, and transparency with a quality similar to that of NeRF-based models. Furthermore, the experiments conducted show that Viewing Direction Gaussian Splatting gives superior results than NeRF and the GS model.